9. a) A manufacture of rayon wants to compare that the yield strength of 5 11⋅ kg/ 2 mm is

met or not at 5% level of significance. The manufacturer draws a sample and

calculates the mean to be 8 12⋅ kg/ 2 mm and the standard derivation is known to be

2⋅0kg/ mm .

2

Carry out the statistical test appropriate for this.


1
Expert's answer
2021-05-24T15:04:26-0400

In this question it is important to test both the null and alternative hypothesis.

"H_o\u200b\\::\u03bc=5.11\\:kg\/mm^2"

"H_1\u200b\\::\u03bc \\neq 5.11\\:kg\/mm^2"

This is a two-tailed test in which the sample standard deviation would be used in a t-test with one mean with unknown population standard deviation. This is a two-tailed test in which the sample standard deviation would be used in a t-test with one mean with unknown population standard deviation.


The significance level is based on the information given as "\u03b1=0.05" and "df=n−1=30−1=29\:\:"degrees of freedom.

Two-tailed test "t_c" critical value is "2.04523."

Region of rejection of this two-tailed test is "R={t:\u2223t\u2223>2.04523}"


"t" - statistic is computed as follows:

"t=\frac{\overline{x}-\mu }{\frac{s}{\sqrt{n}}}=\frac{8.12−5.11}{\frac{2.0}{\sqrt{30}}}=8.24322\approx 8.243"


Therefore, it is inferred that "\:|t|=8.24322>2.04523=|t_c|," hence a conclusion that the null hypothesis is rejected.

There is a concrete proof to claim that the population mean "\mu"  is different than "5.11," at the "\alpha=0.05" significance level.


By utilizing the P value strategy, the significance level for the p-value for two-tailed is:

"\alpha=0.05, t=8.24322" and "df=29" degrees of freedom is "p<0.00001,"

and since p<0.00001<0.05=α it is concluded that the null hypothesis is rejected.


As a result, there is sufficient evidence to conclude that the population mean "\mu" is distinct from  "5.11," at the "\alpha=0.05" significance level.